OpenStudy (anonymous):

3x + 2y = 6 4x + y = 1 Solve the system of equations.

OpenStudy (anonymous):

Do the second equation first

OpenStudy (anonymous):

x= -4/5 y=21/5

OpenStudy (anonymous):

Thank you Gary!!!

OpenStudy (anonymous):

no problem